# Code Readability in the Age of Large Language Models: An Industrial Case Study from Atlassian

### v2

@@
Abstract.
Software engineers spend a significant amount of time reading code during the software development process. This trend is amplified by the emergence of large language models (LLMs) that automatically generate code. However, little is known about the readability of the LLM-generated code and whether it is still important from practitioners’ perspectives in this new era. In this paper, we conduct a survey to explore the practitioners’ perspectives on code readability in the age of LLMs and investigate the readability of our LLM-based software development agents framework, HULA, by comparing its generated code with human-written code in real-world scenarios. Overall, the findings underscore that (1) readability remains a critical aspect of software development; (2) the readability of our LLM-generated code is comparable to human-written code, fostering the establishment of appropriate trust and driving the broad adoption of our LLM-powered software development platform.
@@ Abstract.
booktitle: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering (FSE ’25), June 23–27, 2025, Trondheim, Norway
1. Introduction
Code readability has long been a key focus in software engineering research (tashtoush2013impact; coleman2018aesthetics; fakhoury2018effect; mannan2018towards; alawad2019empirical; johnson2019empirical). For example, Minelli et al. found that programmers dedicate approximately 70% of their time to reading and understanding source code (minelli2015know). Similarly, as Guido Van Rossum, the creator of Python, has noted: “Code is read more often than it is written”(pepeight2001). Thus, writing readable code not only facilitates better comprehension of its functionality (johnson2019empirical) but also significantly reduces software maintenance costs.
In the age of large language models (LLMs) for code, these models have demonstrated remarkable results across various software development tasks (zhao2023survey) (e.g., code completion (takerngsaksiri2024syntax; takerngsaksiri2024student), test case generation (takerngsaksiri2024tdd; alagarsamy2024a3test), code review automation (li2022codereviewer; pornprasit2024gpt; thongtanunam2022autotransform; hong2022commentfinder), vulnerability detection and repair (fu2022linevul; fu2024aibughunter; fu2022vulrepair)). Recently, there has been a growing interest in integrating LLMs into modern Integrated Development Environments (IDEs) (e.g., GitHub Copilot (copilot2024)) to assist software engineers in implementing new features, updating patches, improving code quality, fixing bugs, and resolving software development tasks.
At Atlassian, code readability has always been a cornerstone of software development, as it plays a critical role in facilitating software evolution and maintenance. In the age of large language models (LLMs), like those integrated into Jira (a task management platform) (jiraai2024), the importance of code readability is amplified. Readable code ensures that teams can efficiently collaborate, debug, and enhance their software, reducing technical debt and long-term costs. Recently, we introduced HULA, a human-in-the-loop software development agents framework powered by large language models (takerngsaksiri2024human). While HULA aims to streamline development tasks, a crucial question arises: does readability retain its value in this LLM-assisted paradigm, and if so, why? Additionally, how readable is the code it generates compared to human-written code? Understanding and ensuring code readability in LLM-generated code is vital for Atlassian, since it directly impacts the productivity and satisfaction of software development teams, enabling teams to work efficiently.
@@ 1. Introduction
Figure 1. The user interface of our HULA (Human-in-the-Loop Software Development Agent) framework, seamlessly integrated into Atlassian JIRA (takerngsaksiri2024human).
@@ 2. RQ2)
    Through an examination of the readability of code generated by our LLM-powered code generation (HULA (takerngsaksiri2024human), powered by GPT-4 (openai2023gpt4)) and human-written code using our 144 internal Jira software development tasks, we found that the readability of LLM-generated code is comparable to human-written code (with a negligible to small effect size difference) based on various measurements (e.g., lines of code, code complexity, comment to code ratio). These findings demonstrate that the LLM-generated code can closely resemble human-written code in terms of readability.
@@
Paper Organization. Section LABEL:sec_background describes the background of HULA, our LLM-based software development agents framework. Section 2 presents the motivation and related work. Section 3 and Section 4 present the approach and results of RQ1 and RQ2, respectively. Section 5 discloses the threats to the validity of our study, while Section 6 draws the conclusion.
2. Motivation and Related Works
@@ 2.1. Motivation
Code readability is a cornerstone of software development at Atlassian, particularly as we integrate LLMs into Jira and other platforms (jiraai2024; takerngsaksiri2024human). With the adoption of LLM-powered tools, the size and complexity of codebases are expected to grow rapidly as the majority of the codebase could be generated by LLMs (adam2004). This paradigm shift raises concerns among Atlassian software engineers and customers regarding the readability of LLM-generated code, which may increase technical debts, incur maintenance costs, and negatively affect enterprise coding standards. Therefore, it is of utmost importance to understand whether readability retains its value in this LLM-assisted paradigm, and if so, why? And, how readable is LLM-generated code compared to human-written code?
2.2. Related Works
@@
Code readability is a property that influences how easily a given piece of code can be read and understood (buse2008metric). Over the past 20 years (tashtoush2013impact; coleman2018aesthetics; fakhoury2018effect; mannan2018towards; alawad2019empirical; johnson2019empirical), prior studies have focused on investigating factors influencing code readability and developing metrics, models, and tools to quantify code readability. For example, Buse and Weimer (buse2008metric; buse2009learning) found that the average lines of code and average number of identifiers per line are closely associated with code readability. Similarly, Posnett et al. (posnett2011simpler) identified associations between readability and metrics such as the number of lines, Halstead Volume, and character entropy. Furthermore, Scalabrino et al. (scalabrino2016improving; scalabrino2018comprehensive) emphasized that readability is influenced not only by code structure but also by textual features such as code comments. Moreover, Alawad et al. (alawad2019empirical) observed a negative correlation between readability and code complexity, reporting that higher complexity is often associated with lower readability.
@@
In the age of large language models (LLMs), studying code readability is more crucial than ever. While LLMs have now been seamlessly integrated into software development tools, workflows, and platforms like GitHub Copilot (copilot2024), IntelliJ IDEA (intellij2025), ChatGPT (openai2023gpt4), and Atlassian’s Jira AI (jiraai2024; takerngsaksiri2024human), they also introduce new challenges. Practitioners often raised concerns about the quality of LLM-generated code. For example, Liu et al. (liu2024refining) found that 47% of ChatGPT-generated code snippets suffer from maintainability issues. Majdinasab et al. (majdinasab2024assessing) discovered that 27% of code suggested by GitHub Copilot contains code vulnerabilities. Similarly, Perry et al. (perry2023users) found that users who had access to an AI assistant wrote significantly less secure code than those without access. Such concerns often lead to a lack of trust in the adoption of LLM-powered software development tools.
@@
While many studies flag issues on LLM-generated code in the security aspect, it is still unclear on the readability aspect of how readable is LLM-generated code compared to human-written code. Recently, Madi (al2022readable) found that code generated by GitHub Copilot is comparable in complexity and readability to code written by human pair programmers in the live coding of a controlled environment. However, little is known about the code readability of LLM-generated code on production in real-world scenarios and how the practitioners perceived code readability in the age of large language models. The importance of code readability and the gap in literature lead us to formulate the following research questions:
@@
3. (RQ1) How and why code readability is important in the age of large language models for code?
3.1. Approach
@@
Table 1. Survey questions (excluding demographics).
@@
Survey. We design the 15-minute Google Form with four main sections: the importance of code readability, the challenges of code readability, the state of code readability, and the readability of LLM-generated code. The themes for the Likert scales of agreement questions (e.g., motivations for code readability) are derived from the literature review. Specifically, we categorized factors identified in the literature into themes, which formed the design of our survey questions (i.e., Q1.3, Q2.2, Q3.1). Excluding the demographic questions, the survey questions are shown in Table 1. The lists of derived themes are shown in Table 2 (Q1.3 and Q2.2) and Table 3 (Q3.1). Lastly, the survey obtained Ethical Permission from the Monash University Human Research Ethics Committee (MUHREC, Project ID 42299) before conducting the research.
Table 2. The Importance of Code Readability
@@
No. Description Source 1 Bugs Preventing (piantadosi2022evolution; scalabrino2018comprehensive) 2 Code Complexity (fakhoury2019improving; tashtoush2023notional; alawad2019empirical; holst2021importance; fakhoury2018effect) 3 Code Comprehension (buse2009learning; posnett2011simpler; dorn2012general; scalabrino2018comprehensive; piantadosi2022evolution; sampaio2016software; fakhoury2018effect; johnson2019empirical) 4 Code Maintenance and Evolution (buse2008metric; buse2009learning; dorn2012general; scalabrino2016improving; scalabrino2018comprehensive; posnett2011simpler; sedano2016code; karanikiotis2020data; piantadosi2022evolution; sampaio2016software; lee2013study; piantadosi2020does; johnson2019empirical) 5 Code Standard & Aesthetic (lee2013study; borstler2016beauty; coleman2018aesthetics) 6 Design Quality (scalabrino2018comprehensive; fakhoury2019improving; mannan2018towards; tashtoush2013impact; piantadosi2022evolution; sampaio2016software) 7 Team Collaboration (holst2021importance; sampaio2016software; lee2013study; piantadosi2020does; johnson2019empirical)
Table 3. Coding Factors Related to Code Readability
@@
No. Description Source 1 Code length (e.g., length of the identifier, line, function, file, etc.) (buse2008metric; buse2009learning; dorn2012general; tashtoush2013impact; posnett2011simpler) 2 Code structure/format (e.g., indentation, use of brackets, blank lines, etc.) (buse2008metric; buse2009learning; tashtoush2013impact; dorn2012general) 3 Code style/standard (e.g., PEP8, ESLint, Linter, etc.) (lee2013study; borstler2016beauty; coleman2018aesthetics; fakhoury2019improving) 4 Comment and documentation (buse2008metric; buse2009learning; dorn2012general; tashtoush2013impact; scalabrino2018comprehensive; scalabrino2016improving; karanikiotis2020data) 5 Descriptive naming variable and function (i.e., lexicon and semantics) (tashtoush2013impact; dorn2012general; scalabrino2018comprehensive; scalabrino2016improving; fakhoury2018effect; sedano2016code) 6 Function/file complexity (johnson2019empirical; fakhoury2018effect; tashtoush2023notional; alawad2019empirical; posnett2011simpler; sedano2016code; karanikiotis2020data)
@@
3.2. Results
Table 4 to 7 show the results of this research question, covering four themes of code readability based on 118 practitioners’ responses.
The Importance: 81% of practitioners agree that code readability is important and the top motivation to improve code readability is to reduce maintenance cost and effort in the long term. Table 4 shows that 65% of practitioners answered spending more than half of their time reading code in their professional practice, while 81% of them were aware of the importance of code readability. Particularly, the top three motivations to prioritize code readability are to reduce maintenance costs and effort in the long term, to help the debugging and troubleshooting process, and to improve code comprehension for oneself and others.
Table 4. (RQ1) The Importance of Code Readability from Practitioners’ Perspective.
@@ 3.2. Results
Table 5. (RQ1) The Challenges of Code Readability from Practitioners’ Perspective.
@@
The Challenges: 79% of practitioners agree that the top challenging factor preventing them from prioritizing code readability is time constraints due to project deadlines. Table 5 shows that 70% of practitioners answered they always or often prioritize code readability in their professional practice. However, the top three factors that prevent them from prioritizing the improvement of code readability are the time constraints of project deadlines, the complexity of the codebase or legacy code, and the focus on code functionality over code readability.
Figure 2. The Demographic of Participants in RQ1 ($n$=118).
Table 6. (RQ1) The State of Code Readability from Practitioners’ Perspective.
@@ F. Function/file complexity 82%      5%
The State of Practice: While manually improving code readability is still a dominant method with 68% of practitioners selection, 72% of practitioners agree to consider adopting LLMs to improve code readability. Table 6 shows that the most popular method for improving code readability in professional practice is still done manually by either peer review (41%) or self-improvement (27%). However, 17% of the practitioners have started to adopt LLMs to improve the code readability. The majority of them indicated that they used ChatGPT (n=12) as a tool. In fact, 72% of the practitioners are open to adopting LLMs to improve code. Additionally, the top three factors that practitioners consider when improving code readability are the descriptive naming variables and functions, the comments and documentation, and the complexity of functions or files.
Table 7. (RQ1) The Readability of LLM-generated Code from Practitioners’ Perspective.
@@
The Readability of LLM-generated Code: 39% of practitioners answer they perceive the LLM-generated code as more readable than the human-written code. Table 7 shows that 54% of practitioners answered using LLMs to generate or complete code more than half of their time in professional practice, indicating the current trend of high usage in LLM-generated code. Meanwhile, 39% of practitioners answered they perceive the LLM-generated code as more readable than the human-written code, followed by 34% of them answered both are similarly readable. The top justifications from the follow-up open-text question are logic complexity (n=25), code length (n=15), a number of loop conditions (n=12) and code structure/format (n=12). Specifically, a practitioner who voted LLM-generated code as more readable indicated that “Human code has too many logic conditions to check”. On the other hand, a practitioner who voted Human code as more readable indicates that “I find it easier to follow along.” Interestingly, some practitioners reason that “Actually, Human code is easier to understand in terms of solution, but having a lot of nested loop and conditions makes it harder to read and debug.”, indicates that understanding code and reading code can be as well on the different scales of decision.
4. (RQ2) How readable is the LLM-generated code compared to human-written code in practice?
4.1. Approach

### v3

@@
Abstract
Software engineers spend a significant amount of time reading code during the software development process, especially in the age of large language models (LLMs) that can automatically generate code. However, little is known about the readability of the LLM-generated code and whether it is still important from practitioners’ perspectives in this new era. In this paper, we conduct a survey to explore the practitioners’ perspectives on code readability in the age of LLMs and investigate the readability of our LLM-based software development agents framework, HULA, by comparing its generated code with human-written code in real-world scenarios. Overall, the findings underscore that (1) readability remains a critical aspect of software development; (2) the readability of our LLM-generated code is comparable to human-written code, fostering the establishment of appropriate trust and driving the broad adoption of our LLM-powered software development platform.
Index Terms:
@@ Abstract
I Introduction
Code readability has long been a key focus in software engineering research [1, 2, 3, 4, 5, 6]. For example, Minelli et al. found that programmers dedicate approximately 70% of their time to reading and understanding source code [7]. Similarly, as Guido Van Rossum, the creator of Python, has noted: “Code is read more often than it is written”[8]. Thus, writing readable code not only facilitates better comprehension of its functionality [6] but also significantly reduces software maintenance costs.
In the age of large language models (LLMs) for code, these models have demonstrated remarkable results across various software development tasks [9] (e.g., code completion [10, 11], test case generation [12, 13], code review automation [14, 15, 16, 17], vulnerability detection and repair [18, 19, 20]). Recently, there has been a growing interest in integrating LLMs into modern Integrated Development Environments (IDEs) (e.g., GitHub Copilot [21]) to assist software engineers in implementing new features, updating patches, improving code quality, fixing bugs, and resolving software development tasks.
At Atlassian, code readability has always been a cornerstone of software development, as it plays a critical role in facilitating software evolution and maintenance. In the age of large language models (LLMs), like those integrated into Jira (a task management platform) [22], the importance of code readability is amplified. Readable code ensures that teams can efficiently collaborate, debug, and enhance their software, reducing technical debt and long-term costs. Recently, we introduced HULA, a human-in-the-loop software development agents framework powered by large language models [23]. While HULA aims to streamline development tasks, a crucial question arises: does readability retain its value in this LLM-assisted paradigm, and if so, why? Additionally, how readable is the code it generates compared to human-written code? Understanding and ensuring code readability in LLM-generated code is vital for Atlassian, since it directly impacts the productivity and satisfaction of software development teams, enabling teams to work efficiently.
@@ I Introduction
Figure 1: The user interface of our HULA (Human-in-the-Loop Software Development Agent) framework, seamlessly integrated into Atlassian JIRA [23].
@@ 2. RQ2)
    Through an examination of the readability of code generated by our LLM-powered code generation (HULA [23], powered by GPT-4 [24]) and human-written code using our 144 internal Jira software development tasks, we found that the readability of LLM-generated code is comparable to human-written code (with a negligible to small effect size difference) based on various measurements (e.g., lines of code, code complexity, comment to code ratio). These findings demonstrate that the LLM-generated code can closely resemble human-written code in terms of readability.
@@
Paper Organization. Section II describes the background of HULA, our LLM-based software development agents framework. Section III presents the motivation and related work. Section IV and Section V present the approach and results of RQ1 and RQ2, respectively. Section VI discloses the threats to the validity of our study, while Section VII draws the conclusion.
II Background
Figure 2: An Overview of HULA (Human-in-the-Loop Software Development Agents) [25], seamlessly integrated into Atlassian JIRA Software.
In this section, we briefly describe our Human-in-the-loop LLM-based software development agents framework (HULA) [23]. It is worth noting that in this study, all the human feedback and guidance are discarded to promote the fair comparison between human-written code and HULA-generated code. Figure 1 shows the user interface of HULA seamlessly embedded in Atlassian JIRA Software.
II-A HULA: Human-in-the-Loop Software Development Agents
At Atlassian, we prioritize human expertise to guide and govern AI agents. Rather than pursuing full automation of software development tasks, we designed an LLM-based software development agent to work alongside practitioners as a collaborative assistant. This approach fosters Human-AI synergy [26, 27], where practitioners and AI agents jointly contribute to task completion, ensuring a collaborative software engineering workflow.
Below, we present the overview process of our HULA framework shown in Figure 2. More information can be found in our official announcement [^1] https://www.atlassian.com/blog/atlassian-engineering/hula-blog-autodev-paper-human-in-the-loop-software-development-agents[25] and previous publication [23].
II-A1 Set Context
The workflow starts when a software engineer selects a development task (i.e., a JIRA issue) along with an associated code repository. The description of a JIRA issue typically includes key information to help humans understand the task to be done, such as user stories, definitions of done, acceptance criteria, or example code.
II-A2 Planning
The AI coding agent starts to identify the relevant files in the given repository and generates a coding plan that outlines the necessary modifications. These include a list of relevant files and the specific changes required to address the work item. The software engineer reviews this plan and may either edit the generated plan directly or provide feedback to the AI coding agent for refinement and regeneration. Once the plan satisfies the engineer’s criteria, approval is given to proceed to the implementation phase.
II-A3 Coding
The AI coding agent generates the code and iteratively integrates feedback from validation tools such as compilers and linters until the code meets the required standards (i.e., passes the validation). Then, the software engineer is accountable for reviewing the code, making necessary modifications, or providing feedback for further refinement. Once the generated code is approved, the software engineer may proceed to submit a pull request.
II-A4 Pull Request
Upon approval, the code changes are assembled into a pull request on Bitbucket for review, adhering to the team’s standard workflow. Alternatively, a new branch can be created from the generated code to allow for additional modifications.
In our previous publication [23], we conducted an in-depth evaluation of functional correctness and practitioner acceptance using a multi-stage evaluation framework. However, aspects of code quality – particularly code readability, which directly impacts the maintainability of the code [28, 29, 30, 31] – remained largely unexamined. In this paper, we focus on a detailed assessment of code quality beyond user acceptance (e.g., the merge ratio of pull requests).
III Motivation and Related Works
@@ III-A Motivation
Code readability is a cornerstone of software development at Atlassian, particularly as we integrate LLMs into Jira and other platforms [22, 23]. With the adoption of LLM-powered tools, the size and complexity of codebases are expected to grow rapidly as the majority of the codebase could be generated by LLMs [32]. This paradigm shift raises concerns among Atlassian software engineers and customers regarding the readability of LLM-generated code, which may increase technical debts, incur maintenance costs, and negatively affect enterprise coding standards. Therefore, it is of utmost importance to understand whether readability retains its value in this LLM-assisted paradigm, and if so, why? And, how readable is LLM-generated code compared to human-written code?
III-B Related Works
@@
Code readability is a property that influences how easily a given piece of code can be read and understood [33]. Over the past 20 years [1, 2, 3, 4, 5, 6], prior studies have focused on investigating factors influencing code readability and developing metrics, models, and tools to quantify code readability. For example, Buse and Weimer [33, 28] found that the average lines of code and average number of identifiers per line are closely associated with code readability. Similarly, Posnett et al. [34] identified associations between readability and metrics such as the number of lines, Halstead Volume, and character entropy. Furthermore, Scalabrino et al. [30, 35] emphasized that readability is influenced not only by code structure but also by textual features such as code comments. Moreover, Alawad et al. [5] observed a negative correlation between readability and code complexity, reporting that higher complexity is often associated with lower readability.
@@
In the age of large language models (LLMs), studying code readability is more crucial than ever. While LLMs have now been seamlessly integrated into software development tools, workflows, and platforms like GitHub Copilot [21], IntelliJ IDEA [36], ChatGPT [24], and Atlassian’s Jira AI [22, 23], they also introduce new challenges. Practitioners often raised concerns about the quality of LLM-generated code. For example, Liu et al. [37] found that 47% of ChatGPT-generated code snippets suffer from maintainability issues. Majdinasab et al. [38] discovered that 27% of code suggested by GitHub Copilot contains code vulnerabilities. Similarly, Perry et al. [39] found that users who had access to an AI assistant wrote significantly less secure code than those without access. Such concerns often lead to a lack of trust in the adoption of LLM-powered software development tools.
@@
While many studies flag issues on LLM-generated code in the security aspect, it is still unclear on the readability aspect of how readable is LLM-generated code compared to human-written code. Recently, Madi [40] found that code generated by GitHub Copilot is comparable in complexity and readability to code written by human pair programmers in the live coding of a controlled environment. However, little is known about the code readability of LLM-generated code on production in real-world scenarios and how the practitioners perceived code readability in the age of large language models. The importance of code readability and the gap in literature lead us to formulate the following research questions:
@@
IV (RQ1) How and why code readability is important in the age of large language models for code?
IV-A Approach
@@
TABLE I: Survey questions (excluding demographics).
@@
Survey. We design the 15-minute Google Form with four main sections: the importance of code readability, the challenges of code readability, the state of code readability, and the readability of LLM-generated code. The themes for the Likert scales of agreement questions (e.g., motivations for code readability) are derived from the literature review. Specifically, we categorized factors identified in the literature into themes, which formed the design of our survey questions (i.e., Q1.3, Q2.2, Q3.1). Excluding the demographic questions, the survey questions are shown in Table I. The lists of derived themes are shown in Table II (Q1.3 and Q2.2) and Table III (Q3.1). Lastly, the survey obtained Ethical Permission from the Monash University Human Research Ethics Committee (MUHREC, Project ID 42299) before conducting the research.
TABLE II: The Importance of Code Readability
@@
No. Description Source 1 Bugs Preventing [31, 35] 2 Code Complexity [41, 42, 5, 43, 3] 3 Code Comprehension [28, 34, 29, 35, 31, 44, 3, 6] 4 Code Maintenance and Evolution [33, 28, 29, 30, 35, 34, 45, 46, 31, 44, 47, 48, 6] 5 Code Standard & Aesthetic [47, 49, 2] 6 Design Quality [35, 41, 4, 1, 31, 44] 7 Team Collaboration [43, 44, 47, 48, 6]
TABLE III: Coding Factors Related to Code Readability
@@
No. Description Source 1 Code length (e.g., length of the identifier, line, function, file, etc.) [33, 28, 29, 1, 34] 2 Code structure/format (e.g., indentation, use of brackets, blank lines, etc.) [33, 28, 1, 29] 3 Code style/standard (e.g., PEP8, ESLint, Linter, etc.) [47, 49, 2, 41] 4 Comment and documentation [33, 28, 29, 1, 35, 30, 46] 5 Descriptive naming variable and function (i.e., lexicon and semantics) [1, 29, 35, 30, 3, 45] 6 Function/file complexity [6, 3, 42, 5, 34, 45, 46]
@@
IV-B Results
Table IV to VII show the results of this research question, covering four themes of code readability based on 118 practitioners’ responses.
The Importance: 81% of practitioners agree that code readability is important and the top motivation to improve code readability is to reduce maintenance cost and effort in the long term. Table IV shows that 65% of practitioners answered spending more than half of their time reading code in their professional practice, while 81% of them were aware of the importance of code readability. Particularly, the top three motivations to prioritize code readability are to reduce maintenance costs and effort in the long term, to help the debugging and troubleshooting process, and to improve code comprehension for oneself and others.
TABLE IV: (RQ1) The Importance of Code Readability from Practitioners’ Perspective.
@@ IV-B Results
TABLE V: (RQ1) The Challenges of Code Readability from Practitioners’ Perspective.
@@
The Challenges: 79% of practitioners agree that the top challenging factor preventing them from prioritizing code readability is time constraints due to project deadlines. Table V shows that 70% of practitioners answered they always or often prioritize code readability in their professional practice. However, the top three factors that prevent them from prioritizing the improvement of code readability are the time constraints of project deadlines, the complexity of the codebase or legacy code, and the focus on code functionality over code readability.
Figure 3: The Demographic of Participants in RQ1 ($n$=118).
TABLE VI: (RQ1) The State of Code Readability from Practitioners’ Perspective.
@@ F. Function/file complexity 82%      5%
The State of Practice: While manually improving code readability is still a dominant method with 68% of practitioners selection, 72% of practitioners agree to consider adopting LLMs to improve code readability. Table VI shows that the most popular method for improving code readability in professional practice is still done manually by either peer review (41%) or self-improvement (27%). However, 17% of the practitioners have started to adopt LLMs to improve the code readability. The most commonly indicated tool is ChatGPT (n=12). Although the majority of the practitioners currently improve code manually, 72% of them are open to adopting LLMs to assist with code readability improvement. Additionally, the top three factors that practitioners consider when improving code readability are the descriptive naming variables and functions, the comments and documentation, and the complexity of functions or files.
TABLE VII: (RQ1) The Readability of LLM-generated Code from Practitioners’ Perspective.
@@
The Readability of LLM-generated Code: 39% of practitioners answer they perceive the LLM-generated code as more readable than the human-written code. Table VII shows that 54% of practitioners answered using LLMs to generate or complete code more than half of their time in professional practice, indicating the current trend of high usage in LLM-generated code. Meanwhile, 39% of practitioners answered they perceive the LLM-generated code as more readable than the human-written code, followed by 34% of them answered both are similarly readable. The top justifications from the follow-up open-text question are logic complexity (n=25), code length (n=15), a number of loop conditions (n=12) and code structure/format (n=12). Specifically, a practitioner who voted LLM-generated code as more readable indicated that “Human code has too many logic conditions to check”. On the other hand, a practitioner who voted Human code as more readable indicates that “I find it easier to follow along.” Interestingly, some practitioners reason that “Actually, Human code is easier to understand in terms of solution, but having a lot of nested loop and conditions makes it harder to read and debug.”, indicates that understanding code and reading code can be as well on the different scales of decision.
V (RQ2) How readable is the LLM-generated code compared to human-written code in practice?
V-A Approach